# A classification result for eternal mean convex flows of finite total curvature type

### 18 Mar 2024 v1

@@ 1. Introduction
By convergence to a catenoid “from the outside” we mean specifically that $\lim\limits_{t\to-\infty}M_{t}$ is a catenoid $C$ and that for all times $M_{t}$ lays in the complement of the solid cylinder $C$ bounds. The convergence will be uniform in smooth topology (with multiplicity one) on compact domains. The finite entropy assumption is equivalent to uniform polynomial volume growth and in the context of the mean curvature flow is a fairly ubiquitous background assumption – here in particular we mean its bounded independent of time. By [26], indeed any type of assumption on total curvature along $M_{t}$ that ensures its of finite total curvature for $t$ sufficiently negative (such as a uniform bound, for instance) implies it is of “finite total curvature type” as described above – minimality is not needed in giving condition (a).
As is well known, in the mean curvature flow mean convexity is a very powerful assumption and under the stronger condition of $\alpha$–noncollapsedness such eternal, hence ancient, flows must be convex, and for these there are a number of very strong classification results (for instance [3]). Of course we do not assume noncollapsedness in this work. The finite total curvature type assumption simplifies the topology of the flows in question as well as aids in applying compactness results and understanding the associated limit. Such an eternal mean curvature flow out of a catenoid as indicated above, dubbed the “reapernoid,” was constructed earlier by the author in joint work with A. Payne [22] for all $n\geq 2$. Indeed, considering the construction in Theorem 1.3 of [22] the mean convexity of the eternal solution is obvious and the finite total curvature type assumption above also follows easily. Note that the statement above, unfortunately, does not say anything about the uniqueness of the flow out of the catenoid and imaginably there could be several; note under some symmetry conditions in [22] we prove some partial uniqueness statements in the same spirit as the present one, but here we are in a more general case. Nor does it say anything about the geometry of $M_{t}$ in the case it is homeomorphic to the plane. These issues and more are discussed in the concluding remarks below.
@@
Now we give a sketch of the argument, where throughout $M_{t}$ satisfies the assumptions of the theorem. First, one can see that the topology of $M_{t}$ is a punctured sphere. To see this, denoting the region $M_{t}$ flows into by $H_{t}$, which is well defined by mean convexity, we first oberseve that $\pi_{1}(H_{t}^{c})$ is trivial by White’s theory for mean convex flows [27], namely because it says that the forward time limit of the flow is either empty or a collection of disjoint flat planes. As a consequence, we then show that $H_{t}^{c}$ is given by a ball with some number of solid half cylinders glued on it. If there is only one solid half cylinder attached then $M_{t}$ is topologically a plane, although of course there may be more than one such cylinder attached.
@@
2. Preliminaries
A (smooth) mean curvature flow of hypersurfaces in ${\mathbb{R}}^{n+1}$ is given by a manifold $M^{n}$ and a map $F:M\times I\to{\mathbb{R}}^{n+1}$ satisfying
@@ 2. Preliminaries
where $I\subseteq{\mathbb{R}}$ is some interval. A good introduction to the mean curvature flow can be found in [20]. In some parts of the argument we’ll also refer to its geometric measure theory analogue, the Brakke flow, and an introduction to that can be found in [25]. By a small abuse of notation one often denotes the image of $F$ at a fixed time $t$ by $M_{t}$. A flow is eternal if $I={\mathbb{R}}$, and ancient when at least $I=(-\infty,T]$ for some $T>-\infty$. Under fairly loose hypotheses the mean curvature flow exists for at least a short time from a given initial smooth hypersurface, for instance compactness or bounded curvature, but producing ancient (or eternal) mean curvature flows is more delicate one reason being that singularities occur often along the flow. One natural way they appear is in fact in the analysis of these singularities, and one would expect to naturally encounter mean convex ancient or eternal flows as singularities along a mean convex mean curvature flow– although oftentimes (such as for compact mean convex initial data) such flows will be noncollapsed so the singularities will in fact be convex. See [9, 10, 3] for more about the noncollapsing assumption and its consequences. The (nontrivial) eternal singularity models one finds will be bowl solitons which are topologically planar and in particular none flow out of a catenoid. On the other hand as mentioned above the mean curvature flow is, as well known, the natural analogue of the heat equation so in this sense ancient and eternal flows, as analogues of complete solutions to the heat equation, are arguably interesting regardless of whatever convexity condition or soliton equation they may satisfy or not.
@@
For later reference, that singularities occur often is a consequence of the comparison principle, which says that two disjoint hypersurface where at least one of them is compact must stay disjoint under the mean curvature flow. In particular, in comparison with the bowl soliton there can be no compact eternal flows.
@@
one sees by the maximum principle it is typically preserved – at least in the compact case and for noncompact ones under some extra assumptions as well (in our setting we are just assuming mean convexity at all times). The mean convex mean curvature flow has been considered by many authors. Particularly important in our applications is the work of B. White, where one can see from sections 10, 11 of his paper [27] the following:
@@
Of course since we are considering minimal surfaces in ${\mathbb{R}}^{3}$ one can further conclude the convergence above will be smooth, although it may be with multiplicity. In particular, the one sided minimization property of mean convex mean curvature flows gives that the limiting set, if nonempty, will be smooth and stable within any bounded domain which says the limiting set in totality is stable – it and some other facts from [27] will play an important role in the following section.
A useful way to characterize mean convex flows is in terms of the corresponding evolution of the sets they bound – because $M_{t}$ is mean convex it bounds a set $H_{t}$ which, in terms of set inclusion, is monotonically decreasing as $t$ increases. Strictly so, because $H$ is assumed to be strictly positive. Because $M_{t}$ is assumed connected it and its complement will be connected as well. We will refer to this characterization often in the arguments below.
@@ 3. Proof of theorem
By White’s theorem from the section above we know the limit set, which we’ll denote here by $M_{\infty}$, will be either empty or a (possibly disconnected) smooth and stable minimal surface – considering each component at a time without loss of generality it is conencted. If it is empty we are finished so suppose $M_{\infty}\neq\emptyset$. Since the flow is mean convex we see that $P\subset H_{t}$ for all $t\in{\mathbb{R}}$, and because $M_{t}$ is connected then the flow must lay on one side of $P$. The multiplicty bound theorem of White [27] then says that $M_{t}$ must converge to $P$ with multiplicity one. Since $M_{t}$ embedded $M_{\infty}$ is as well, and it will be closed so complete. Using it is stable and complete then by [6] it must be a plane $P$. $\square$
@@
For the sake of concreteness we consider the fate of the flow starting from time $t=0$: because the flow is smooth for any $s,t\in{\mathbb{R}}$ we have $H_{s}^{c}$ and $H_{t}^{c}$ are homotopic. We recall that $H_{t}$ denotes the set $M_{t}$ bounds and flows monotonically into. Considering then a homotopically nontrivial curve $\gamma\in H_{0}^{c}$, note by the mean convexity of the flow that $H_{t}^{c}\subset H_{s}^{c}$ for $t<s$ so for $t>0$ we may sensically consider $\gamma\subset H_{t}^{c}$. Since the flow is smooth, clearly $\gamma$ must not be homotopically trivial in $H_{t}^{c}$ for any finite $t>0$. On the other hand by lemma 3.1 it must be the case that $\gamma$ is homotopically trivial in $H_{\infty}^{c}$, or so that it bounds an immersed disc in $H_{\infty}^{c}$. Of course the disc lays in a compact set, so because by lemma 3.1 $H_{t}^{c}$ converges to $H_{\infty}^{c}$ which is uniform in compact sets after perhaps perturbing the disc slightly it must be contained in $H_{t}^{c}$ for some potentially large but finite $t$ though, giving a contradiction because the flow is smooth. $\square$
@@
$H_{0}^{c}$ is homeomorphic to a ball with a finite number of solid half cylinders glued along it so in particular, referring to part (a) of the defintion of finite total curvature type, $\widetilde{M}$ must be a sphere.
@@
Suppose that the genus $g$ of $\widetilde{M}$ is greater than or equal to 1. Again fixing the time $t=0$ to be concrete, by capping off the ends of $M_{0}$ into the solid half cylinders of $H_{0}^{c}$ using the disjoint discs from the lemma above and smoothing out the edges we obtain a smooth surface $\Sigma$, homeomorphic to $\widetilde{M}$, which bounds a domain $U$. Because the capped ends are all homeomorphic to planes, one can see by Seifert–Van Kampen and that ${\mathbb{R}}^{3}$ is simply connected that the domains in $H_{0}^{c}$ they bound are simply connected. Using this one can continuously deform any closed curve in $H_{0}^{c}$ into one laying entirely $U$ so that $\pi_{1}(U)\simeq\pi_{1}(H_{0}^{c})$. Since $\overline{U}$ is a compact 3–manifold, we have that the rank of the image of the map $H^{1}(\overline{U})\to H_{1}(\Sigma)$ is half the rank of $H_{1}(\Sigma)$ (see lemma 3.5 of [11]). When $g\geq 1$ the rank of $H_{1}(\Sigma)$ is nonzero of course, which implies that $H^{1}(\overline{U})$ and hence $H_{1}(\overline{U})$ is nontrivial, as a consequence of the universal coefficient theorem. So, by the Hurwicz isomorophism, neither is $\pi_{1}(U)$ contradicting lemma 3.2. $\square$
To refine this picture, we will glean information about $M_{t}$ from its limit in the far past i.e. as $t\to-\infty$. As above first we examine the geometry of the limit. The convergence as well as the geometry will be improved in the lemma afterwards.
@@
The set $\lim\limits_{t\to-\infty}M_{t}$ is either empty or diffeomorphic to a disjoint union of smooth embedded complete minimal surfaces, where the convergence is in the sense of varifolds possibly with multiplicity.
@@
For the sake of exposition first let’s give an argument in the case $M_{t}$ has pointwise uniformly bounded curvature, by which we mean $|A|^{2}(p,t)<C$ for some fixed constant $C$ independent of $p$ and $t$– this is indeed satisfied by the eternal flow out of a catenoid mentioned in the introduction, but apriori this isn’t immediate. Note that by the pointwise curvature bound and Shi’s estimates that $|\nabla^{k}A|$ for any $k\geq 1$ is bounded uniformly along $M_{t}$ for all times $t<0$. This gives by the entropy bound and Arzela–Ascoli subsequential convergence of the flows $M_{t+s_{i}}$ for a sequence $s_{i}\to-\infty$ in $C^{k}$ norm (apriori, possibly multisheeted) for any $k\geq 0$ on bounded sets. The limit flow will be smooth (but potentially only immersed) and evolve by the classical mean curvature flow. Naturally we’ll denote this limit flow, which for all we know at this point depends on $s_{i}$, by $M_{-\infty,t}$.

### 14 Mar 2026 v2

@@ 1. Introduction
By convergence to a catenoid “from the outside” we mean specifically that $\lim\limits_{t\to-\infty}M_{t}$ is a catenoid $C$ and that for all times $M_{t}$ lays in the complement of the solid cylinder $C$ bounds. The convergence will be uniform in smooth topology (with multiplicity one) on compact domains. The finite entropy assumption is equivalent to uniform polynomial volume growth and in the context of the mean curvature flow is a fairly ubiquitous background assumption – here in particular we mean it’s bounded independent of time. By [27], indeed any type of assumption on total curvature along $M_{t}$ that ensures it’s of finite total curvature for $t$ sufficiently negative (such as a uniform bound, for instance) implies it is of “finite total curvature type” as described above – minimality is not needed in giving condition (a).
As is well known, in the mean curvature flow mean convexity is a very powerful assumption and under the stronger condition of $\alpha$–noncollapsedness such eternal, hence ancient, flows must be convex, and for these there are a number of very strong classification results (for instance [3]). Of course we do not assume noncollapsedness in this work. The finite total curvature type assumption simplifies the topology of the flows in question as well as aids in applying compactness results and understanding the associated limit. Such an eternal mean curvature flow out of a catenoid as indicated above, dubbed the “reapernoid,” was constructed earlier in [22] for all $n\geq 2$. Indeed, considering the construction in Theorem 1.3 of [22] the mean convexity of the eternal solution is obvious and the finite total curvature type assumption above also follows easily. Note that the statement above, unfortunately, does not say anything about the uniqueness of the flow out of the catenoid and imaginably there could be several; note under some symmetry conditions in [22] we prove some partial uniqueness statements in the same spirit as the present one, but here we are in a more general case. Nor does it say anything about the geometry of $M_{t}$ in the case it is homeomorphic to the plane. These issues and more are discussed in the concluding remarks below.
@@
Now we give a sketch of the argument, where throughout $M_{t}$ satisfies the assumptions of the theorem. First, one can see that the topology of $M_{t}$ is a punctured sphere. To see this, denoting the region $M_{t}$ flows into by $H_{t}$, which is well defined by mean convexity, we first obrserve that $\pi_{1}(H_{t}^{c})$ is trivial by White’s theory for mean convex flows [28], namely because it says that the forward time limit of the flow is either empty or a collection of disjoint flat planes. As a consequence, we then show that $H_{t}^{c}$ is given by a ball with some number of solid half cylinders glued on it. If there is only one solid half cylinder attached then $M_{t}$ is topologically a plane, although of course there may be more than one such cylinder attached.
@@
2. Preliminaries on the mean curvature flow
In this section we give a nonexhaustive account of some of the results and notions in this paper we will use below, primarily focusing on those concerning the flow. We start with the most basic definition:
Definition 2.1.
A (smooth) mean curvature flow of embedded hypersurfaces in ${\mathbb{R}}^{n+1}$ is given by a manifold $M^{n}$ and a family of embeddings $F:M\times I\to{\mathbb{R}}^{n+1}$ satisfying
@@ 2. Preliminaries on the mean curvature flow
where $I\subseteq{\mathbb{R}}$ is some nonempty interval.
By a small abuse of notation one often denotes the image of $F$ at a fixed time $t$ by $M_{t}$. In the following we discuss some especially pertinent properties of the flow, a more complete introduction can be found in [20].
A flow is eternal if $I={\mathbb{R}}$, and ancient when at least $I=(-\infty,T]$ for some $T>$-$\infty$. Under fairly loose hypotheses the mean curvature flow exists for at least a short time from a given initial smooth hypersurface, for instance compactness or uniformly bounded curvature, but producing ancient (or eternal) mean curvature flows is more delicate one reason being that singularities occur often along the flow.
@@
One natural way they appear is in the analysis of these singularities; we say that a flow $M_{t}$ develops a singularity at a point $(p,t)$ when $|A|^{2}(p_{i},t_{i})\to\infty$ for a sequence $(p_{i},t_{i})\to(p,T)$. To study these singularities one parabolically rescales the flow along such a sequence by factors $\lambda_{i}\to 0$ and, when appropriate weak assumptions are met to employ compactness theorems are satisfied (see the discussion of Brakke flows below, for example) one will find in a subsequential limit an ancient, possibly eternal, limiting flow which can be thought of as models of the singularity. Of course, one would expect to naturally encounter mean convex ancient or eternal flows as models of singularities along a mean convex mean curvature flow – although oftentimes (such as for compact mean convex initial data) such flows will be noncollapsed so the singularities will in fact be convex. See [9, 10, 3] for more about the noncollapsing assumption and its consequences. The (nontrivial) eternal singularity models one finds in these cases will be bowl solitons which are topologically planar and in particular none flow out of a catenoid. On the other hand as mentioned above the mean curvature flow is, as well known, the natural analogue of the heat equation so in this sense ancient and eternal flows, as analogues of complete solutions to the heat equation, are arguably interesting regardless of whatever convexity condition or soliton equation they may satisfy or not.
@@
It is a consequence of the comparison principle, which says that two disjoint hypersurfaces where at least one of them is compact must stay disjoint under the mean curvature flow, that singularities occur often. In particular, in comparison with the bowl soliton there can be no compact eternal flows. Due to the ubiquity of singularities weak notions of the flow are crucial in many situations, and there are a number of ways to define them. A particularly important notion of weak flow is the Brakke flow, a geometric measure theoretic definition which we’ll refer to at times in the argument below. These are defined as follows:
Definition 2.2.
A (n-dimensional integral) Brakke flow is a family of Radon measures $\mu_{t}$ such that:
  1. (1)
[…]
  2. (2)
    For a bounded interval $t_{1},t_{2}]\subset I$ and any compact set $K$,
    $\int_{t_{1}}^{t_{2}}\int_{K}(1+H^{2})d\mu_{t}dt<\infty$ (2.2)
  3. (3)
    (Brakke inequality) For all compactly supported nonnegative test functions $\phi$,
    $\int_{V(t_{0})}\phi d\mu_{t_{0}}\leq\int_{V(0)}\phi d\mu_{0}+\int_{0}^{t_{0}}\int_{V(t)}\phi H^{2}-H\langle\nabla\phi,\nu\rangle-\frac{d\phi}{dt}d\mu_{t}dt$ (2.3)
The study of Brakke flows itself is quite rich, a comprehensive introduction to Brakke flows can be found in [26], but let’s briefly discuss a couple facts which will be needed about them. First is Brakke’s compactness theorem, which says that for a sequence of Brakke flows with uniform area bounds on parabolic cylinders one may exact a converging subsequence. A second important fact is Brakke’s regularity theorem, which says that for Brakke flows with density bounds in a backwards parabolic neighborhood sufficiently close to 1 will be smooth with bounded curvature in a smaller neighborhood. There are related facts for varifolds, Allard compactness and Allard regularity respectively, under analogous assumptions which we will also refer to; see [24]. Although in this paper we only consider smooth flows, because our assumptions generally aren’t strong enough to employ, for instance, Arzela–Ascoli compactness we need to consider these wider classes of objects.
The entropy $\lambda(M_{t})$, introduced by Colding and Minicozzi in [5], of a flow $M_{t}$ is defined as the supremum over recentered and rescaled Gaussian functionals:
$\lambda(M_{t})=\sup\limits_{x\in{\mathbb{R}}^{3},r\in(0,\infty)}(4\pi r)^{-n/2}\int_{\Sigma}e^{-|x-x_{0}|/4r}d\mu(x)$ (2.4)
This is a quantity that has been of intense recent interest and has many useful properties, but the only point in stipulating the flows $M_{t}$ in theorem 1.1 above have bounded entropy is that it provides a compact means of writing that $M_{t}$ has time independent area bounds to employ the compactness results mentioned above and is a commonly made background assumption.
@@
one sees by the maximum principle it is typically preserved – at least in the compact case and for noncompact ones under some extra assumptions as well (in our setting we are just assuming mean convexity at all times). The mean convex mean curvature flow has been considered by many authors. Particularly important in our applications is the work of B. White [28]. We will use a number of his results from this article, a result one can see from sections 10, 11 of his paper we wish to highlight is the following:
@@
Of course since we are considering minimal surfaces in ${\mathbb{R}}^{3}$ one can further conclude the convergence above will be smooth, although it may be with multiplicity. In particular, the one sided minimization property of mean convex mean curvature flows gives that the limiting set, if nonempty, will be smooth and stable within any bounded domain which says the limiting set in totality is stable.
Lastly, a useful way to characterize mean convex flows is in terms of the corresponding evolution of the sets they bound – because $M_{t}$ is mean convex it bounds a set $H_{t}$ which, in terms of set inclusion, is monotonically decreasing as $t$ increases. Strictly so in fact, because $H$ is assumed to be strictly positive. Because $M_{t}$ is assumed connected it and its complement will be connected as well. We will refer to this characterization often in the arguments below.
@@ 3. Proof of theorem
By White’s theorem, theorem 2.1, from the section above we know the limit set, which we’ll denote here by $M_{\infty}$, will be either empty or a (possibly disconnected) smooth and stable minimal surface – considering each component at a time without loss of generality it is connected. If it is empty we are finished so suppose $M_{\infty}\neq\emptyset$. Since the flow is mean convex we see that $P\subset H_{t}$ for all $t\in{\mathbb{R}}$, and because $M_{t}$ is connected then the flow must lay on one side of $P$. The multiplicity bound theorem of White [28] then says that $M_{t}$ must converge to $P$ with multiplicity one. Since $M_{t}$ embedded $M_{\infty}$ is as well, and it will be closed so complete. Using it is stable and complete then by [6] it must be a plane $P$. $\square$
@@
For the sake of concreteness we consider the fate of the flow starting from time $t=0$: because the flow is smooth for any $s,t\in{\mathbb{R}}$ we have $H_{s}^{c}$ and $H_{t}^{c}$ are homotopic. We recall that $H_{t}$ denotes the set $M_{t}$ bounds and flows monotonically into. Considering then a homotopically nontrivial curve $\gamma\in H_{0}^{c}$, note by the mean convexity of the flow that $H_{t}^{c}\subset H_{s}^{c}$ for $t<s$ so for $t>0$ it makes sense to consider $\gamma\subset H_{t}^{c}$. Since the flow is smooth, clearly $\gamma$ must not be homotopically trivial in $H_{t}^{c}$ for any finite $t>0$. On the other hand by lemma 3.1 it must be the case that $\gamma$ is homotopically trivial in $H_{\infty}^{c}$, or so that it bounds an immersed disc in $H_{\infty}^{c}$. Of course the disc lays in a compact set, so because by lemma 3.1 $H_{t}^{c}$ converges to $H_{\infty}^{c}$ which is uniform in compact sets after perhaps perturbing the disc slightly it must be contained in $H_{t}^{c}$ for some potentially large but finite $t$ though, giving a contradiction because the flow is smooth. $\square$
@@
$H_{0}^{c}$ is homeomorphic to a ball with a finite number of solid half cylinders glued along it so in particular, referring to part (a) of the definition of finite total curvature type, $\widetilde{M}$ must be a sphere.
@@
Suppose that the genus $g$ of $\widetilde{M}$ is greater than or equal to 1. Again fixing the time $t=0$ to be concrete, by capping off the ends of $M_{0}$ into the solid half cylinders of $H_{0}^{c}$ using the disjoint discs from the lemma above and smoothing out the edges we obtain a smooth surface $\Sigma$, homeomorphic to $\widetilde{M}$, which bounds a domain $U$. Because the capped ends are all homeomorphic to planes, one can see by Seifert–Van Kampen (using slightly enlarged domains bounded by these capped off ends) and that ${\mathbb{R}}^{3}$ is simply connected that the domains in $H_{0}^{c}$ they bound are indeed simply connected. Using this one can continuously deform any closed curve in $H_{0}^{c}$ into one laying entirely $U$ so that $\pi_{1}(U)\simeq\pi_{1}(H_{0}^{c})$. Since $\overline{U}$ is a compact 3–manifold, we have that the rank of the image of the map $H^{1}(\overline{U})\to H_{1}(\Sigma)$ is half the rank of $H_{1}(\Sigma)$ (see lemma 3.5 of [11]). When $g\geq 1$ the rank of $H_{1}(\Sigma)$ is nonzero of course, which implies that $H^{1}(\overline{U})$ and hence $H_{1}(\overline{U})$ is nontrivial, as a consequence of the universal coefficient theorem. So, by the Hurwicz isomorophism, neither is $\pi_{1}(U)$ contradicting lemma 3.2. $\square$
To refine this picture, we will glean information about $M_{t}$ from its limit in the far past i.e. as $t\to$-$\infty$. The first lemma says that the limit is well defined/the convergence is full, at least in a weak sense, essentially because the sets $H_{t}$ are monotone increasing in terms of set inclusion as $t\to$-$\infty$:
@@
The limits $\lim_{t\to-\infty}H_{t}$, $\lim_{t\to-\infty}M_{t}$, which we denote by $_{-\infty}, H\partial H_{-\infty}$ respectively, are well defined/the convergence is full in the Hausdorff topology on bounded domains.